# Branch actions and the structure lattice

### v1

@@ 4.3. Proof of A
as $\mathrm{Supp}(\mathrm{rist}_{G}(v)^{(k)})$ is contained in the subtree rooted at $v$. Thus $\mathrm{rist}_{G}(v)^{(k)}$ has the same number of distinct conjugates as $\mathrm{rist}_{G}(v)$, namely the number of vertices in the $G$-orbit of $v$ by the orbit-stabilizer theorem. By the level-transtivity of $G$ this is precisely the number of vertices at the same level as $v$.
Now by Proposition 4.2, the support of $H$ is clopen and a disjoint union of cone sets. Let $\mathrm{Supp}(H)=\bigcup_{i=1}^{n}C_{v_{n}}$. Then the above reasoning applies to each $v_{i}$ and we get
$H\geq\prod_{i=1}^{n}\mathrm{rist}_{G}(v_{i})^{(k_{i})}$
for some $k_{i}\geq 0$. We may assume $v_{1},\dotsc,v_{n}$ are ordered by graded lexicographical order. Let us also consider $k\geq 1$ such that $k\geq k_{i}$ for every $1\leq i\leq n$. Then
(4.4) $H\geq\big{(}\mathrm{rist}_{G}(v_{n})^{H}\big{)}^{(k)}$
as $\mathrm{rist}_{G}(v_{n})^{H}$ is a direct product and the derived subgroups of a direct product are no more than the direct product of the derived subgroups of the factors. Now
$\big{(}\mathrm{rist}_{G}(v_{n})^{H}\big{)}^{(k)}\leq_{\mathrm{va}}\big{(}\mathrm{rist}_{G}(v_{n})^{H}\big{)}^{(k-1)}\leq_{\mathrm{va}}\dotsb\leq_{\mathrm{va}}\mathrm{rist}_{G}(v_{n})^{H}$
@@
(4.5) $\Big{[}\big{(}\mathrm{rist}_{G}(v_{n})^{H}\big{)}^{(k)}\Big{]}=\Big{[}\big{(}\mathrm{rist}_{G}(v_{n})^{H}\big{)}^{(k-1)}\Big{]}=\dotsb=\Big{[}\mathrm{rist}_{G}(v_{n})^{H}\Big{]}$
@@
$H\cap\mathrm{rist}_{G}(v_{n})^{H}\leq_{\mathrm{va}}\mathrm{rist}_{G}(v_{n})^{H}.$
Now $\mathrm{rist}_{G}(v_{n})^{H}$ has finite index in
$\mathrm{Stab}_{G}(\partial T\setminus\mathrm{Supp}(H))\leq G.$
Since $H\leq\mathrm{Stab}_{G}(\partial T\setminus\mathrm{Supp}(H))$, we also get
$H\cap\mathrm{rist}_{G}(v_{n})^{H}\leq_{f}H.$
@@
$[H]=\Big{[}\mathrm{rist}_{G}(v_{n})^{H}\Big{]}=\Big{[}\prod_{v\in\mathrm{Supp}(H)}\mathrm{rist}_{G}(v)\Big{]}$
@@
$\mathrm{rist}_{G}(v_{n})^{H}$ $\leq_{f}\prod_{v\in\mathrm{Supp}(H)}\mathrm{rist}_{G}(v).\qed$

### v2

@@ 4.3. Proof of A
as $\mathrm{Supp}(\mathrm{rist}_{G}(v)^{(k)})$ is contained in $C_{v}$. Thus $\mathrm{rist}_{G}(v)^{(k)}$ has the same number of distinct conjugates as $\mathrm{rist}_{G}(v)$, namely the number of vertices in the $G$-orbit of $v$ by the orbit-stabilizer theorem. By the level-transitivity of $G$ this is precisely the number of vertices at the same level as $v$.
Now by Proposition 4.3, the support of $H$ is clopen and a disjoint union of cone sets. Let $\mathrm{Supp}(H)=\bigsqcup_{i=1}^{n}C_{v_{i}}$, where we assume as before that each $v_{i}$ is moved by $H$. Then the above reasoning applies to each $v_{i}$ and we get
$H\geq\prod_{i=1}^{n}\mathrm{rist}_{G}(v_{i})^{(k)}.$
We may assume $v_{1},\dotsc,v_{n}$ are all at the same level by replacing if necessary a vertex with all its descendants at a lower level. Then
(4.4) $H\geq\prod_{i=1}^{n}\mathrm{rist}_{G}(v_{i})^{(k)}=\big{(}\prod_{i=1}^{n}\mathrm{rist}_{G}(v_{i})\big{)}^{(k)}$
as the derived subgroup of a direct product is no more than the direct product of the derived subgroups of the factors. Now
$\big{(}\prod_{i=1}^{n}\mathrm{rist}_{G}(v_{i})\big{)}^{(k)}\leq_{\mathrm{va}}\big{(}\prod_{i=1}^{n}\mathrm{rist}_{G}(v_{i})\big{)}^{(k-1)}\leq_{\mathrm{va}}\dotsb\leq_{\mathrm{va}}\prod_{i=1}^{n}\mathrm{rist}_{G}(v_{i})$
@@
(4.5) $\Big{[}\big{(}\prod_{i=1}^{n}\mathrm{rist}_{G}(v_{i})\big{)}^{(k)}\Big{]}=\Big{[}\big{(}\prod_{i=1}^{n}\mathrm{rist}_{G}(v_{i})\big{)}^{(k-1)}\Big{]}=\dotsb=\Big{[}\prod_{i=1}^{n}\mathrm{rist}_{G}(v_{i})\Big{]}$
@@
$H\cap\prod_{i=1}^{n}\mathrm{rist}_{G}(v_{i})\leq_{\mathrm{va}}\prod_{i=1}^{n}\mathrm{rist}_{G}(v_{i}).$
Now we have
$\prod_{i=1}^{n}\mathrm{rist}_{G}(v_{i})\leq_{f}\mathrm{Stab}_{G}(\partial T\setminus\mathrm{Supp}(H))\leq G$
as $G$ is branch and
$\prod_{i=1}^{n}\mathrm{rist}_{G}(v_{i})=\mathrm{Rist}_{G}(N)\cap\mathrm{Stab}_{G}(\partial T\setminus\mathrm{Supp}(H)),$
where $N$ is the common level of $T$ at which all the vertices $v_{1},\dotsc,v_{n}$ lie. Since $H\leq\mathrm{Stab}_{G}(\partial T\setminus\mathrm{Supp}(H))$, we also get
$H\cap\prod_{i=1}^{n}\mathrm{rist}_{G}(v_{i})\leq_{f}H.$
@@
$[H]=\Big{[}\prod_{i=1}^{n}\mathrm{rist}_{G}(v_{i})\Big{]}=\Big{[}\prod_{v\in\mathrm{Supp}(H)}\mathrm{rist}_{G}(v)\Big{]}$
@@
$\prod_{i=1}^{n}\mathrm{rist}_{G}(v_{i})=\mathrm{Rist}_{G}(N)\cap\prod_{v\in\mathrm{Supp}(H)}\mathrm{rist}_{G}(v)$ $\leq_{f}\prod_{v\in\mathrm{Supp}(H)}\mathrm{rist}_{G}(v).\qed$